A family's monthly income is $2,800 and they spend $700each month on food. Write the amount they spend on food as a fraction of their monthly income in lowest terms.
Answer by Theo(13342)

their monthly income is 2800
their monthly food bill is 700.

the amount they spend on food as a fraction of their monthly food is 700/2800 = 7/28 = 1/4

their monthly income is 2800
1/4 * 2800 = 2800/4 = 700